OpenStudy (anonymous):

somebody help me with this equation f(x)=x^2+4x<0

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

can you solve it for me

OpenStudy (anonymous):

oh.. ok. so you can treat the "<" as an "=" in these types of problems, then just factor out an x to get: x(x+4)=0 then set x=0 and x+4=0 and just solve for x and you will get your two answers

OpenStudy (anonymous):

show me how plz

OpenStudy (anonymous):

x^2+4x=0 x(x+4)=0 x=0 and x+4=0 so x=0 & -4

OpenStudy (anonymous):

whats the square root of x+4,x>equal to 0

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

can you solve that problem for me

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

if i scan my work in send it would you help me with all them'

OpenStudy (anonymous):

i cant. sorry but i have to study for my own test in about 5 min

OpenStudy (anonymous):

can u help me with my math plz